OpenStudy (anonymous):

A rock is thrown into a still pond. The circular ripples move outward from the point of impact of the rock so that the radius of the circle formed by a ripple increases at the rate of 5 feet per minute. Find the rate at which the area is changing at the instant the radius is 17 feet.

OpenStudy (anonymous):

\[ A=\pi r^2 \]

OpenStudy (anonymous):

\[ \frac{dA}{dt}=\frac{dA}{dr}\frac{dr}{dt}=(2\pi r)\frac{dr}{dt} \]

OpenStudy (anonymous):

How about at 7ft radius?

OpenStudy (anonymous):

" The circular ripples move outward from the point of impact of the rock so that the radius of the circle formed by a ripple increases at the rate of 5 feet per minute." This means \(dr/dt=5\)

OpenStudy (anonymous):

"the radius is 17 feet" \(r=17\)

OpenStudy (anonymous):

"Find the rate at which the area is changing" Find \(dA/dt\)

OpenStudy (anonymous):

\[ \frac{dA}{dt}=(2\pi r)\frac{dr}{dt}=2\pi (17)(5) \]

OpenStudy (anonymous):

@Sarah1362894 Use the algebra. Let go.

OpenStudy (anonymous):

I try and I'm wrong everytime

OpenStudy (anonymous):

The answer is \(170\pi\) isn't it?

OpenStudy (anonymous):

Nope just tried it

OpenStudy (anonymous):

did you put in the unit of ft squared per minute?